OpenStudy (anonymous):

show all work to solve (1/x) + (3/8) =(1/2) explain how to solve this equation

OpenStudy (anonymous):

subtract \[\frac{3}{8}\] from both sides take the reciprocal of both side

Directrix (directrix):

1/x + 3/8 = 1/2 Multiply through by 8 8/x + 3 = 4 8/x = 1 8/x= 1/1 x = 8